Lisa can shovel the drive in 35 minutes. Bill can do the same in 40 minutes.How long would it take if the shoveled the drive together
Answer by unlockmath(1688)

Hello,
The equation we want to set up is:
1/35+1/40=1/x Multiply by 1400x to get:
40x+35x=1400 Combine like terms to get:
75x=1400 Divide 75 into both sides to get:
x=18 2/3 minutes
There we go.
RJ
www.math=unlock.com